OpenStudy (anonymous):

for the sequence given by a1=3 and an=2an-1-4 find the first three terms

OpenStudy (sidsiddhartha):

\[a_1=3\\a_n=2a_{n-1}-4\] that it?

OpenStudy (sidsiddhartha):

first put n=2 u'll get\[a_2=2a_{2-1}-4=2a_1-4=(2 \times 3)-4=6-4=2\] next:put n=3

OpenStudy (sidsiddhartha):

having trouble?

OpenStudy (anonymous):

then what sidsiddhartha

OpenStudy (sidsiddhartha):

then u'll get\[a_3=2a_2-4=(2 \times 2)-4=0\\then~again~put,n=4\\a_4=2a_3-4=(2 \times 0)-4=-4\] so next three terms are=2,0,-4

OpenStudy (anonymous):

thank you so much :)

OpenStudy (sidsiddhartha):

yw!! :)

OpenStudy (anonymous):

do you think you would have time to help me with one more problem

OpenStudy (sidsiddhartha):

yeah i can :)

OpenStudy (anonymous):

are you good with graphs

OpenStudy (sidsiddhartha):

i'll try

OpenStudy (anonymous):

Im not sure how to write the problem on here